# MaskMoE: Boosting Token-Level Learning via Routing Mask in Mixture-of-Experts

### v2

@@ 4.1 Pre-training Dataset
Following previous works, we use the Pile dataset Gao et al. (2021) as pre-training data. The Pile dataset is a large-scale, publicly available corpus, containing 22 domains and over 825GB of English text data. For our experiments, we use the well-known LLaMA tokenizer for tokenization, with a vocabulary size of $32k$. We follow  Xie et al. (2023a); Su et al. (2024) to calculate the sampling rate for each domain based on the number of tokens after tokenization. Due to the computational budget, we follow the pretraining settings of  Xie et al. (2023a); Su et al. (2024); Huang et al. (2024); Xiong et al. (2024); Lian et al. (2024), all models are pre-trained with 100B tokens.
Then, to identify infrequent tokens and frequent tokens, we calculate the frequency of each token in the training set and sort them in descending order of frequency. We categorize the top tokens that cover $P\times 100\%$ of the dataset as frequent ones, and the remaining $(1-P)\times 100\%$ as infrequent ones, where $P$ is a tunable hyperparameter, ranging from 0 to 1.
@@ 4.2 Compared Models
We compare the proposed MaskMoE with four models for validation experiments.
@@
  • SMoE denotes a Switch Transformer Fedus et al. (2022), where the router is a learnable parameter. Unless otherwise specified, each MoE layer has 64 experts.
  • Hash Layer Roller et al. (2021) signifies that through a random hash method, each token is assigned a fixed expert prior to training. Unless otherwise specified, each MoE layer has 64 experts.
  • Share-MoE is a hybrid dense and MoE model created using residual connections Rajbhandari et al. (2022). Models with shared experts are currently a popular architecture Dai et al. (2024); DeepSeek-AI et al. (2024); Zhao et al. (2024); Rajbhandari et al. (2022). Unless otherwise specified, Share-MoE has 1 shared expert and 128 routed experts, where each expert is 0.5 times the size of a standard FFN. During both training and inference, in addition to the activation of shared experts, the top expert is also selected from a pool of 128 experts. In such a setup, Share-MoE has the same number of floating-point operations (FLOPs) and activation parameters as SMoE and Hash Layer. This allows for a fair comparison between Share-MoE, SMoE, and Hash Layer.
  • MaskMoE aims to achieve a balance between representation diversity in MoE layers and thorough parameter training, via the proposed routing masking method. Similarly to Share-MoE, MaskMoE also employs the remarkable shared expert architecture, which has been proven by  Dai et al. (2024); Rajbhandari et al. (2022) to be beneficial for the training of MoE models.
@@ 4.3 Experimental Setup
Following  Touvron et al. (2023a, b); Huang et al. (2024), we adopt the LLaMA architecture with 24 transformer layers and a hidden layer dimension of 1024 and we employ the AdamW  Loshchilov and Hutter (2019) optimizer with a cosine learning rate decay schedule.
For the MoE model, following  Roller et al. (2021), we conduct experiments under both single-layer and multi-layer settings. For the single-layer setting, we replace the final FFN sublayer with the MoE sublayer. For the multi-layer MoE models, we follow Gshard Lepikhin et al. (2021), and the FFN sublayer with MoE sublayer for MoE layers every other layer, resulting in a total of $12$ MoE layers in this setting. For single-layer MoE models and dense models, we use a learning rate of $3e^{-4}$, while for $12$-layer MoE models, we employ a learning rate of $1e^{-4}$ to ensure stable convergence of the models.
If not specified otherwise, tokens that account for the top 40% of frequency are considered frequent tokens, while the remaining are considered infrequent tokens, i.e., $P=0.4$. Empirically, for frequent tokens, the visible expert count $V$ is set to 8, and for infrequent tokens, the visible expert count $V$ is set to 1. All of our implementations are based on the DeepSpeed [^1] https://github.com/microsoft/DeepSpeed library Rajbhandari et al. (2022); Rasley et al. (2020), which offers robust support for MoE-distributed training. By default, we enable the random token selection method Kim et al. (2021) implemented in the library to facilitate faster model convergence and better runtime efficiency. Following previous works Roller et al. (2021); Fedus et al. (2022); Dai et al. (2022), unless otherwise specified, our experiments only select the top expert.
@@ 4.4 Main Results
In this section, we first present the model’s PPL on the Pile validation set. Then, following Touvron et al. (2023a); Brown et al. (2020); Su et al. (2024); Dai et al. (2024), we conduct tests on a large number of downstream benchmarks, including zero-shot tests for BoolQ Clark et al. (2019), HellaSwag Zellers et al. (2019), LAMBADA Paperno et al. (2016), PIQA Bisk et al. (2020), SIQA Sap et al. (2019), StoryCloze Mostafazadeh et al. (2016), and Arc-e Bhakthavatsalam et al. (2021), as well as 5-shot tests for TriviaQA Joshi et al. (2017) and WebQs Berant et al. (2013). Among them, TriviaQA and WebQs use exact match as the metric, while the remaining benchmarks are evaluated based on accuracy. For a fair comparison, we use the open-source evaluation tool lm-evaluation-harness [^2] https://github.com/EleutherAI/lm-evaluation-harness for assessment.
@@ 4.4.1 Perplexity Results
Table 1 shows the main results of language modeling on the Pile validation sets. With the same number of activated parameters during inference (i.e., 468M), MoE models consistently achieve improvements over the dense model. In comparison to SMoE and Hash Layer, under the Single-MoE-Layer setup, MaskMoE has respectively reduced the PPL on the Pile Validation by 0.14 and 0.08. In the Multi-MoE-Layer setup, their corresponding PPL are, respectively, reduced by 0.07 and 0.05. Compared to Share-MoE, MaskMoE achieves a PPL decrease of 0.24 in Single-MoE-Layer setup and another PPL decrease of 0.04 in Multi-MoE-Layer setup. Such results demonstrate the reasonableness of the MaskMoE design.
It is noteworthy that Share-MoE achieves worse results than SMoE and Hash Layer in the Single-MoE-Layer setup. However, MaskMoE consistently maintains superior performance, whether in single or multi-MoE layers. It indicates that MaskMoE exhibits stronger robustness across various settings.
@@ 4.4.2 Benchmark Results
As shown in Table 2, we present the performance of the models on downstream tasks. Firstly, we observe that MoE models consistently and significantly outperform the dense models in downstream tasks. And more importantly, we find that MaskMoE substantially outperforms the other MoE baselines, whether in Single-MoE-Layer setup or Multi-MoE-Layer setup.
Compared to SMoE, MaskMoE consistently outperforms it in all 9 benchmarks with the Single-MoE-Layer setup and 8 out of 9 benchmarks with the Multi-MoE-Layer setup. Against Hash Layer, MaskMoE excels in 7 out of 9 benchmarks with the Single-MoE-Layer setup and in 8 out of 9 benchmarks with the Multi-MoE-Layer setup. Compared to Share-MoE, MaskMoE demonstrates significant improvements in 7 out of 9 benchmarks with the Single-MoE-Layer setup and in 6 out of 9 benchmarks with the Multi-MoE-Layer setup.
The result above well demonstrates the performance enhancement achieved by using our proposed router masking method to control the visibility of experts for different tokens. We believe the improvements are due to two factors. Firstly, compared to SMoE and Share-MoE, infrequent tokens are only routed to a single expert, ensuring more thorough training for those tokens. Secondly, in contrast to the Hash Layer, frequent tokens have multiple expert options, which promotes diversity in the representation learning of frequent tokens.
@@ 5.1 Impact of Shared Expert
To explore the impact of shared experts, we conduct ablation experiments on them (i.e., removing the shared experts), and the results are shown in Table 3. We separately report the average scores [^3] Due to the limited space, here we simply average all metrics on downstream tasks to get an average score for each model. of the benchmarks and the PPL of Pile Validation. Note that after removing the shared experts, we adopt the same architecture as SMoE and Hash Layer, where the MoE layer consists of 64 fully-sized FFN.
We find that even without shared experts, MaskMoE still outperforms SMoE, Hash Layer, and Share-MoE in average benchmark scores and PPL. Such a result further validates the soundness of our proposed MaskMoE. Moreover, incorporating shared expert structures continuously improves MaskMoE’s performance. We attribute it to that the shared expert structure promotes greater specialization in token representation DeepSeek-AI et al. (2024).
5.2 Impact of Router Mask
Model Avg($\uparrow$) PPL($\downarrow$)
SMoE 43.87 6.62
Hash Layer 44.56 6.56
Share-MoE 44.25 6.72
MaskMoE 44.99 6.48
w/o Shared Experts 44.85 6.51
Table 3: Average scores benchmarks and PPL on the Pile validation sets. The best score is marked in bold, and the second best is underlined.
Considering the proposed MaskMoE without shared experts still outperforms the compared MoE methods, to show the impact of router mask more significantly and minimize the impact of the shared expert, the experiments remove the shared expert. Instead, we adopt the same structure as SMoE and Hash Layer, with each MoE layer comprising 64 full-sized FNN modules as experts.
The performance of MaskMoE is predominantly influenced by two hyperparameters: the boundary threshold for the separation of frequent and infrequent tokens, and the maximum number of visible experts (i.e., $V$) for tokens. As shown in Table 4, we conduct an experimental parameter search for both and report their PPL values on the Pile validation sets. $V_{a}$ represents the number of experts visible for frequent tokens, and $V_{b}$ represents the number of experts visible for infrequent tokens. Notably, when $V_{a}=V_{b}=1$, MaskMoE is equivalent to the Hash Layer, and when $V_{a}=V_{b}=64$, MaskMoE is equivalent to SMoE.
Then we have the following observations. Firstly, we find that MaskMoE outperforms the Hash Layer. With $V_{a}=8$ and $V_{b}=1$, MaskMoE reduces ppl by 0.052 on the Pile validation set compared to the Hash Layer, i.e., $V_{a}=V_{b}=1$. The result indicates that the multiple-expert design in the proposed MaskMoE for frequent tokens does enhance their representation diversity, which is beneficial for model performance.
Secondly, compared to SMoE, concentrating tokens into fewer visible experts, such as settings like $V_{a}=V_{b}=8$ or $V_{a}=V_{b}=4$, yields better outcomes. It indicates that routing tokens more densely to a reduced number of visible experts can enhance thorough training and improve model performance. Moreover, we observe that the configuration $V_{a}=8$ and $V_{b}=1$ outperforms $V_{a}=V_{b}=8$. It well verifies that further reducing the number of visible experts is more beneficial for infrequent tokens.
Finally, it is noteworthy that for MaskMoE, when $V_{b}=1$, regardless of the value of $V_{a}$ (whether it is 4, 8, 16, or 64), the performance of MaskMoE consistently outperforms that of SMoE and Hash Layer. That also demonstrates the robustness of MaskMoE.
@@ 5.2 Impact of Router Mask
0.2 8 1 6.518 0.4 16 1 6.509
0.4 8 1 6.506 0.4 4 1 6.511
0.6 8 1 6.518 0.4 64 1 6.549
/ 64 64 6.618$\spadesuit$ / 1 1 6.558 $\clubsuit$
/ 4 4 6.551 / 8 8 6.566
Table 4: Impact of Router Mask. $\spadesuit$ denotes SMoE, $\clubsuit$ denotes Hash Layer.
5.3 Number of Experts
To investigate the variation in model performance under different configurations of expert numbers, we conducted experiments with 32, 64, and 128 experts. The experimental results, as shown in Figure 2, reveal the following observations.
First of all, we observe that for SMoE and Share-MoE, increasing the number of experts does not necessarily lead to a lower PPL. In fact, when the number of experts reaches 128, both models even exhibit higher PPLs compared to having only 64 experts. We attribute it to that having more experts leads to the occurrences of an identical token being spread among more experts, resulting in each expert not learning enough about this token, which leads to model underfitting. In contrast, the performance of Hash Layer and MaskMoE continue to improve with the increase in the number of experts. We attribute it to that, despite the growing number of experts, the visible experts for an identical token in Hash Layer and MaskMoE remain a fixed, small number, allowing the experts to continue learning adequately. It indicates that they are more favorable for scaling the number of experts.
Secondly, we observe that compared to the Hash Layer, MaskMoE consistently performs better across all configurations. Although the Hash Layer also enables experts to be adequately trained with the tokens assigned to them, MaskMoE excels in the diversity of representations learning for frequent tokens. such a superiority in handling token diversity contributes to MaskMoE’s overall advantage over Hash Layer.
Finally, we find that regardless of the number of experts, MaskMoE consistently maintains the best performance. That further verifies the superiority of the proposed MaskMoE and the reasonableness of its design.

### v4

@@ 4.1 Pre-training Dataset
Following previous works, we use the Pile dataset Gao et al. (2021) as pre-training data. The Pile dataset is a large-scale, publicly available corpus, containing 22 domains and over 825GB of English text data. For our experiments, we use the well-known LLaMA tokenizer for tokenization, with a vocabulary size of $32k$. We follow  Xie et al. (2023a); Su et al. (2024) to calculate the sampling rate for each domain based on the number of tokens after tokenization. Due to the limited computational budget, and following the pretraining settings of  Xie et al. (2023a); Su et al. (2024); Huang et al. (2024); Xiong et al. (2024); Lian et al. (2024), all models are pre-trained with 100B tokens.
Then, to identify infrequent tokens and frequent tokens, we calculate the frequency of each token in the training set and sort them in descending order of frequency. We categorize the top tokens that cover $P\times 100\%$ of the dataset as frequent ones, and the remaining as infrequent ones, where $P$ is a tunable hyperparameter, ranging from 0 to 1.
@@ 4.2 Compared Models
We compare the proposed MaskMoE with four remarkable baselines for validation experiments. Following previous works Roller et al. (2021); Fedus et al. (2022); Dai et al. (2022), unless otherwise specified, our experiments select the top-$1$ expert for each MoE layer.
@@
  • SMoE denotes a Switch Transformer Fedus et al. (2022), where the router is a learnable layer. Unless otherwise specified, each MoE layer has 64 experts.
  • Hash Layer Roller et al. (2021) signifies that through a random hash method, each token is assigned to a fixed expert before training. Unless otherwise specified, each MoE layer has 64 experts.
  • Share-MoE is a hybrid dense and MoE model created using residual connections Rajbhandari et al. (2022). Models with shared experts are currently a popular architecture Dai et al. (2024); DeepSeek-AI et al. (2024); Zhao et al. (2024); Rajbhandari et al. (2022). Unless otherwise specified, Share-MoE has 1 shared expert and 128 routed experts, where each expert is 0.5$\times$ of the size of a standard FFN. During both training and inference, in addition to the activation of shared experts, the top-$1$ expert is also selected from the pool of 128 experts. In such a setup, Share-MoE has the same number of floating-point operations (FLOPs) and activation parameters as SMoE and Hash Layer. That allows for a fair comparison between Share-MoE, SMoE, and Hash Layer.
  • MaskMoE aims to achieve a balance between representation diversity and thorough training, via the proposed routing masking method. Similarly to Share-MoE, MaskMoE also employs the remarkable shared expert architecture, which has been proven by  Dai et al. (2024); Rajbhandari et al. (2022) to be beneficial for the training of MoE models.
@@ 4.3 Experimental Setup
Following  Touvron et al. (2023a, b); Huang et al. (2024), we adopt the LLaMA architecture with 24 Transformer blocks and a hidden-state dimension of 1024. We employ the AdamW  Loshchilov and Hutter (2019) optimizer with a cosine learning rate decay schedule.
For the MoE model, following  Roller et al. (2021), we conduct experiments under both single-layer and multi-layer settings. For the single-layer setting, we replace the FFN layer with the MoE layer only for the last Transformer block. For the multi-layer MoE models, we follow Gshard Lepikhin et al. (2021), and replace the FFN layer with MoE layer for every other Transformer block, resulting in a total of $12$ MoE layers in this setting. For single-layer MoE models and dense models, we use a learning rate of $3e^{-4}$, while for Multi-layer MoE models, following  Lewis et al. (2021), we employ a learning rate of $1e^{-4}$ to ensure stable convergence of the models.
If not specified otherwise, the top most frequent tokens that cover 40% of the training set are considered frequent tokens, while the remaining are considered infrequent tokens, i.e., $P=0.4$. Empirically, for frequent tokens, the visible expert count $V$ is set to 8, and for infrequent tokens, the visible expert count $V$ is set to 1. All of our implementations are based on the DeepSpeed [^1] https://github.com/microsoft/DeepSpeed library Rajbhandari et al. (2022); Rasley et al. (2020), which offers robust support for MoE-distributed training. By default, we enable the random token selection method Kim et al. (2021) implemented in the library to facilitate faster model convergence and better runtime efficiency.
@@ 4.4 Main Results
In this section, we first present the model’s PPL on the Pile validation set. Then, following Touvron et al. (2023a); Brown et al. (2020); Su et al. (2024); Dai et al. (2024), we conduct tests on various downstream benchmarks, including zero-shot tests for BoolQ Clark et al. (2019), HellaSwag Zellers et al. (2019), LAMBADA Paperno et al. (2016), PIQA Bisk et al. (2020), SIQA Sap et al. (2019), StoryCloze Mostafazadeh et al. (2016), and Arc-e Bhakthavatsalam et al. (2021). Following  Touvron et al. (2023a); Su et al. (2024), we conduct 5-shot tests for TriviaQA Joshi et al. (2017) and WebQs Berant et al. (2013). Among them, TriviaQA and WebQs use exact match as the metric, while the remaining benchmarks are evaluated based on accuracy. For a fair comparison, we use the open-source evaluation tool lm-evaluation-harness [^2] https://github.com/EleutherAI/lm-evaluation-harness for assessment.
@@ 4.4.1 Perplexity Results
Table 1 shows the main results of language modeling on the Pile validation sets. With the same number of activated parameters during inference (i.e., 468M), MoE models consistently achieve improvements (i.e., substantially lower PPL) over the dense model. In comparison to SMoE and Hash Layer, under the Single-MoE-Layer setup, MaskMoE has respectively reduced the PPL on the Pile validation set by 0.14 and 0.08. In the Multi-MoE-Layer setup, their corresponding PPL are, respectively, reduced by 0.07 and 0.05. Compared to Share-MoE, MaskMoE also achieves a PPL decrease of 0.24 in Single-MoE-Layer setup and 0.04 in Multi-MoE-Layer setup. Such results demonstrate the effectiveness of our proposed MaskMoE.
It is noteworthy that Share-MoE achieves worse results than SMoE and Hash Layer in the Single-MoE-Layer setup, even with a built-in shared expert. However, MaskMoE consistently maintains superiority over them in both single-layer and multi-layer setups. That indicates MaskMoE exhibits stronger robustness across various settings.
@@ 4.4.2 Benchmark Results
As shown in Table 2, we present the performance of the models on downstream tasks. We can observe that MoE models consistently and significantly outperform the dense models in downstream tasks. More importantly, MaskMoE substantially outperforms the other MoE baselines, whether in Single-MoE-Layer setup or Multi-MoE-Layer setup.
Specifically, Compared to SMoE, MaskMoE consistently outperforms it in all 9 benchmarks with the Single-MoE-Layer setup and 8 of 9 benchmarks with the Multi-MoE-Layer setup. Against Hash Layer, MaskMoE excels in 7 of 9 benchmarks with the Single-MoE-Layer setup and in 8 of 9 benchmarks with the Multi-MoE-Layer setup. Compared to Share-MoE, MaskMoE demonstrates significant improvements in 7 of 9 benchmarks with the Single-MoE-Layer setup and in 6 of 9 benchmarks with the Multi-MoE-Layer setup.
The result above well demonstrates the performance enhancement achieved by using our proposed router masking method that controls the visibility of experts for different tokens. We believe the improvements are due to two factors. Firstly, compared to SMoE and Share-MoE, infrequent tokens are only routed to a single expert, ensuring more thorough training for those tokens. Secondly, in contrast to the Hash Layer, frequent tokens have multiple expert options, which promotes diversity in their representation learning.
@@ 5.1 Impact of Shared Expert
To explore the impact of shared experts, we conduct ablation experiments on them (i.e., removing the shared experts), and the results are shown in Table 3. We separately report the performance on the benchmark and the PPL on the Pile validation set. Note that after removing the shared experts, we adopt the same architecture as SMoE and Hash Layer, where the MoE layer consists of 64 fully-sized FFN.
We find that even without Shared-Experts, MaskMoE still outperforms SMoE, Hash Layer, and Share-MoE in terms of benchmark scores and PPL, with Table 2 as references. Specifically, compared to SMoE, MaskMoE w/o Shared-Experts excels in 8 of 9 benchmarks. Additionally, compared to Hash Layer and Share-MoE, MaskMoE w/o Shared-Experts gains significant improvements in 7 of 9 benchmarks. Similarly, MaskMoE w/o Shared-Experts, yields lower PPL than SMoE, Hash Layer, and Share-MoE, with Table 1 as references. Such a result further validates the soundness of our proposed MaskMoE. Moreover, incorporating Shared-Expert structures continuously improves MaskMoE’s performance. As shown in Table 3, compared to MaskMoE w/o Shared-Expert, MaskMoE excels in 5 of 9 benchmarks and has a lower PPL in the Pile validation set. We attribute it to that the shared expert structure promotes greater specialization in token representation DeepSeek-AI et al. (2024).
MaskMoE w/o Shared-Expert
BoolQ 58.38 57.40
Hellaswag 44.47 44.10
LAMBADA 55.36 54.86
PIQA 68.99 69.10
SIQA 41.86 41.97
StoryCloze 65.15 65.69
Arc-e 53.14 53.91
TriviaQA 10.39 10.01
WebQs 7.14 6.89
PPL($\downarrow$) 6.48 6.51
Table 3: Impact of Shared Expert on benchmark performances and the Pile validation PPL. The best score is marked in bold.
@@
0.0 8 1 6.558 $\spadesuit$ 0.6 8 1 6.518
0.2 8 1 6.518 0.8 8 1 6.539
0.4 8 1 6.506 1.0 8 1 6.566 $\spadesuit$
Table 4: Impact of the frequency threshold $P$ for the separation of frequent and infrequent tokens. $\spadesuit$ denotes no distinction between frequent and infrequent tokens.
$P$ $V_{a}$ $V_{b}$ PPL $P$ $V_{a}$ $V_{b}$ PPL
/ 64 64 6.618$\spadesuit$ 0.4 64 1 6.549
/ 8 8 6.566 0.4 8 1 6.506
/ 4 4 6.551 0.4 4 1 6.511
/ 1 1 6.558 $\clubsuit$ 0.4 16 1 6.509
Table 5: Impact of visible experts. $\spadesuit$ denotes SMoE, $\clubsuit$ denotes Hash Layer.
5.2 Impact of Hyperparameters for Router Masking
Considering the proposed MaskMoE without shared experts still outperforms the compared MoE methods, to show the impact of router mask more clear and minimize the impact of the shared expert, here we keep the same experimental setting with the shared expert removed and conduct further experiments to analyze the impact of hyperparameters of our proposed MaskMoE.
The performance of MaskMoE is predominantly influenced by two hyperparameters: the boundary threshold (i.e., $P$) for the separation of frequent and infrequent tokens, and the maximum number of visible experts (i.e., $V$) for tokens. As shown in Table 4 and 5, we conduct parameter searches for both and report their PPL values on the Pile validation sets. Specifically, here $V_{a}$ represents the number of experts visible for frequent tokens, and $V_{b}$ represents the number of experts visible for infrequent tokens.
As shown in Table 4, we report the impact of the boundary threshold $P$. Here, $P=0.0,V_{a}=8,V_{b}=1$ and $P=1,V_{a}=8,V_{b}=1$ indicate no distinction between frequent and infrequent tokens, with every token having $1$ and $8$ visible experts, respectively. Their performances are significantly worse than those under the settings of $10<P<$ 1, given that $V_{a}=8,V_{b}=1$. That indicates tokens of different frequencies indeed require distinct visible experts, validating the effectiveness of our design. Additionally, we observe that regardless of the settings of $P$ where $10<P<$ 1, the PPLs are consistently lower than $P=0.0$ or $P=1$. That validates the robustness of MaskMoE.
[…]
@@ 5.2 Impact of Hyperparameters for Router Masking
5.3 Impact of the Number of Experts
To investigate the variation in model performance under different configurations of expert numbers, we conducted experiments with 32, 64, and 128 experts. The experimental results are illustrated in Figure 2.
First of all, we observe that for SMoE and Share-MoE, increasing the number of experts does not necessarily lead to a lower PPL. In fact, when the number of experts reaches 128, both models even exhibit higher PPLs compared to having only 64 experts. We attribute it to that having more experts leads to the occurrences of an identical token being spread among more experts, and that disables each expert to learn adequately about this token, leading to model underfitting. In contrast, the performance of Hash Layer and MaskMoE continue to improve as the number of experts increases. Actually, when the number of experts increases, the average number of token occurrences assigned to each expert decrease. However, Hash Layer and MaskMoE can still ensure that each expert is fed with all or a large portion of the occurrences of a token, though the number of distinct tokens decreases for each expert. That makes the data for each expert more consistent and easier to learn.
Secondly, compared to Hash Layer, MaskMoE consistently performs better across all configurations. Although Hash Layer also enables experts to be adequately trained for specific tokens, MaskMoE excels in the diversity of representation learning for frequent tokens, which contributes to MaskMoE’s overall advantage over Hash Layer.
Finally, regardless of the number of experts, MaskMoE consistently yields the best performance. That further verifies its superiority and reasonableness.
5.4 Performance Analysis of Top-k Routing
Model SMoE Hash Layer Share-MoE MaskMoE
PPL($\downarrow$) 6.55 6.50 6.58 6.46
Table 6: The PPL on the Pile validation sets under the top-$2$ gating. The best score is marked in bold.
We further analyze the performance of the MoE models under the top-$k$ ($k=2$) routing mechanism. In the SMoE model, the top-$2$ experts are selected from 64 fully-sized FFN experts based on the router’s output scores. For the Hash Layer, two experts are randomly chosen from 64 fully-sized FFN experts using a random hash function. Regarding Share-MoE and MaskMoE models, in addition to the shared experts, the top-$2$ experts are selected from the 128 half-sized FFN experts based on the router’s output scores. All settings share the same number of active parameters. Note that in MaskMoE, the number of visible experts is doubled from the original setting, i.e., $V_{a}=16,V_{b}=2$. As shown in Table 6, the model’s PPLs on the Pile validation sets are presented. MaskMoE consistently yields the best performance, achieving a 0.12 decrease in PPL compared to Share-MoE. That further verifies the effectiveness.